For these problems I need use the discriminant to determine the number of solutions of the quadratic equation, and whether the solutions are real or complex. It is not necessary to find the roots; just determine the number and types of solutions.
2x^2 + 5x + 5 = 0

Answer by stanbon(75887)

just determine the number and types of solutions.
2x^2 + 5x + 5 = 0
---
a = 2 ; b = 5 ; c = 5
---
discriminant = b^2-4ac = 25 -4*2*5 = -15
=================================================
Since the discriminant is negative the
quadratic has two unequal Complex Number solutions.
=================================================
Cheers,
Stan H